OpenStudy (anonymous):

HELP! for arbitrary sets A, B, C, and D. (AxB)U(CxD)=(AUC)x(BUD). True or false? Prove your answer.

OpenStudy (zarkon):

what do you think?

OpenStudy (anonymous):

False? I really am not good at understanding sets and set theory

OpenStudy (zarkon):

what would you get if A and D were empty sets?

OpenStudy (anonymous):

BUC?

OpenStudy (anonymous):

=CXB

OpenStudy (zarkon):

if \(A=D=\emptyset\) then \(A\times B=\emptyset\) and \(C\times D=\emptyset\) and therefore \((A\times B)\cup(C\times D)=\emptyset\)

OpenStudy (zarkon):

\(A=D=\emptyset\) then \(A\cup C=C\) and \(B\cup D=B\) thus \((A\cup C)\times(B\cup D)=C\times B\)

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

okay but they are arbitrary set

OpenStudy (anonymous):

which means we don't know if they are empty or not

OpenStudy (zarkon):

I just gave you a blueprint on how to construct a counter example